# MEMO: Dataset and Methods for Robust Multimodal Retinal Image Registration with Large or Small Vessel Density Differences

### v1

@@ 5.4 Evaluation Metrics
5.4.1 RMSE
Based on the predicted registration matrices, the labeled points in all test OCTA images were reprojected to the corresponding test EMA images. Then, the root-mean-square-error (RMSE) [42] between the reprojected points and the labeled points was computed.
@@ 5.4.2 Success rate
The success rate was defined as the number of image pairs with successful registration over the total number of test image pairs. A registration was successful when its RMSE $<$ 10, based on the clinical tolerance.
5.4.3 Soft Dice (SD)/ Masked Soft Dice (MSD)
Dice is widely used to evaluate the registration quality by calculating the pixel alignment between the warped source vessel masks and the target vessel masks. Soft Dice (SD) [27] has been proposed as an extension of Dice for assessing registration quality. For SD, CLAHE [63] was first applied to enhance the contrast of two input images and the Frangi vesselness filter [64] was then used to generate the vesselness probability masks of the two input images. We calculated SD by
$SD=\frac{2\sum_{i=1}^{N}min(F_{i}^{s^{\prime}},F_{i}^{t})}{\sum_{i=1}^{N}F_{i}^{s^{\prime}}+\sum_{i=1}^{N}F_{i}^{t}},$ (8)
@@ 5.4.3 Soft Dice (SD)/ Masked Soft Dice (MSD)
Figure 5: The average (a) SD and (b) MSD values over image pairs in MEMO by adding different x and y shifts to the ground truth registration. The top-left value in each figure represents the average SD or MSD value obtained by ground truth registration. All values are color-coded.
Additionally, we found that SD could not accurately represent performance when a relatively large VD difference was present between the two modalities, such as in our MEMO dataset. Moreover, it was particularly unreliable when the results of each competing method had relatively small differences. As most vessels in OCTA images do not exist in the corresponding EMA images, calculating SD based on every pixel is not ideal. Hence, when evaluating on the MEMO dataset, we extended SD to Masked Soft Dice (MSD) which considered only the pixels within the ground truth segmentation masks of EMA images. MSD is defined as
$MSD=\frac{2\sum_{i=1}^{N}min(M_{i}^{e^{\prime}}F_{i}^{e^{\prime}},M_{i}^{e^{\prime}}F_{i}^{o})}{\sum_{i=1}^{N}M_{i}^{e^{\prime}}F_{i}^{e^{\prime}}+\sum_{i=1}^{N}M_{i}^{e^{\prime}}F_{i}^{o}},$ (9)
where $F^{e^{\prime}}$ and $F^{o}$ are the vesselness probability masks of the warped EMA images and the OCTA images, and $M^{e^{\prime}}$ represents the warped ground truth segmentation masks of EMA images. In Fig. 5, we demonstrated that MSD has better ability to assess the registration performance on our MEMO dataset, as it is less sensitive to the VD difference and noise.
@@ 6.1 The CF-FA Dataset
Methods Success rate RMSE SD
Before Registration 0.00% 68.35 -
SURF-PIIFD-RPM 82.76% 15.21 0.51
LoFTR 41.28% 45.13 0.51
SG 82.76% 23.68 0.52
@@
$*{}^{*}$ 89.67% 14.03 0.51
VDD-Reg (Ours) 100.00% 3.19 0.51
$*{}^{*}$he authors of Content-Adaptive [28] have demonstrated better performance on the CF-FA dataset by fine-tuning SuperPoint and adding an outlier rejection network.
@@ 6.2 The MEMO Dataset
Methods Success rate RMSE SD MSD
Before Registration 6.67% 88.48 - -
SURF-PIIFD-RPM 6.67% 58.48 0.42 0.41
LoFTR 13.33% 53.55 0.43 0.49
SG 0.00% 257.01 0.28 0.44
@@
Content-Adaptive 33.33% 78.69 0.41 0.48
VDD-Reg (Ours) 86.67% 21.99 0.50 0.61
[…]
Table 4 shows the quantitative registration results of our method and the baseline methods on our MEMO dataset. Due to the relatively large VD difference, all the methods performed worse compared to the results on the CF-FA dataset. Still, our VDD-Reg outperformed all the baseline methods by a large margin (86.67%). From Table 4, we observed that SURF-PIIFD-RPM performed poorly (6.67%) on our MEMO dataset, which suggested that the hand-crafted features might be insufficiently powerful for this scenario. The two deep learning-based direct methods, LoFTR and SG, also produced unsatisfactory results (13.33% and 0%) due to the large distribution gap between their training dataset [66] and our MEMO dataset. CycleGAN-based with pretrained SuperPoint (SP) achieved relatively better performance compared to other competing methods, demonstrating its potential to solving difficult multimodal registration problems. Content-Adaptive, which is also a segmentation-based method, performed much worse (33.33%) than our method on the MEMO dataset. We attributed this to our use of annotated vessel segmentation masks from a single modality (EMA in our case). Different from our two-stage semi-supervised learning framework (LVD-Seg), Content-Adaptive trained the segmentation networks naively with style loss, the self-comparison loss and the image registration loss. To improve the segmentation quality, Content-Adaptive additionally guided the segmentation networks with mean phase images of input images using pixel-adaptive convolution (PAC) [51]. However, due to the high complexity of OCTA images, the OCTA mean phase images were usually too noisy to correctly guide the segmentation networks. On the other hand, our LVD-Seg framework used very few (e.g., three) annotated vessel segmentation masks from one modality to guide the segmentation networks to segment similar vessels in EMA and OCTA image pairs for both stages. In addition, the two-stage design also enhanced training stability when style loss was involved. These are particularly important for multimodal retinal image registration when a large VD difference exists between the two modalities.
Figure 6: Registration results of our method and the baseline methods on a selected image pair from our MEMO dataset. The top row shows the grid images where the EMA and OCTA images are interlaced as small grids. The bottom row shows the overlay images of the EMA (green) and OCTA (orange) vessel segmentation masks generated by each method. The RMSE and MSD of each method are listed below each overlay image.
In Figure 6, we also demonstrated the qualitative registration results of our method and the baseline methods on a selected image pair from our MEMO dataset. Two different approaches were used to present the results, including grid images (top) and overlay images (bottom). The RMSE and MSD are listed below the overlay images. Our method demonstrated more accurate alignment compared to all the baseline methods.
@@ 7.1 Ablation Study on the Two-stage Learning Framework
Method Success rate RMSE SD MSD
@@
In this section, we investigated the benefits of each stage in the proposed two-stage semi-supervised learning framework (LVD-Seg) by removing one of the stages from the framework. The results are shown in Table 5. For Stage 1 only, we trained the EMA vessel segmentation network following the procedure described in Section 4.1.2 and used the same EMA vessel segmentation network for segmenting OCTA images. For Stage 2 only, we trained the segmentation networks with style loss only. In general, the performance of both variants decreased significantly. Stage 1 only performed the worst probably due to the limited number (three in our case) of annotated vessel segmentation masks used for supervised training, making the segmentation network generalize poorly on the test images and resulting in poor registration performance. Furthermore, Stage 1 only directly applied the trained EMA segmentation network on OCTA images, which may not work due the relatively large VD difference between the two modalities. Although Stage 2 only worked better than Stage 1 only, it still significantly lagged behind our LVD-Seg. This implied that relying purely on style loss could let the training become unstable and resulted in unreliable registration. All these results emphasized the effectiveness of the proposed two-stage semi-supervised learning framework (LVD-Seg).
@@ 7.2 Ablation Study on Number of Required Vessel Masks
# of annotated masks Success rate RMSE SD MSD
@@ 7.3 Ablation Study on Data Used for Supervised Training
Datasets Success rate RMSE SD MSD
@@ 8 Conclusion
In this paper, we present MEMO, the first public multimodal EMA and OCTA retinal image dataset. MEMO provides registration ground truth, EMA image sequences and OCTA projection images, desirable for various research fields. With MEMO, we first uncover a unique challenge of multimodal retinal image registration between modalities with large VD differences. After that, we propose a segmentation-based deep-learning registration framework, VDD-Reg, and a new evaluation metric, Masked Soft Dice (MSD), to deal with the large vessel density difference between EMA and OCTA in multimodal retinal image registration. Moreover, to train the segmentation module in our VDD-Reg, we design a novel two-stage semi-supervised learning framework, LVD-Seg, which combines supervised and unsupervised losses. Both quantitative and qualitative results demonstrate that VDD-Reg outperforms the baseline methods in both small VD differences (i.e., CF-FA) and large VD differences (i.e., MEMO). Additionally, VDD-Reg requires as few as three annotated vessel segmentation masks to maintain its performance, which demonstrates its promising potential for registering other modalities.
Acknowledgment
The authors would like to acknowledge the funding support from the University of Maryland Strategic Partnership: MPowering the State, a formal collaboration between the University of Maryland College Park and the University of Maryland Baltimore.

### v2

@@ 5.4 Evaluation Metrics
5.4.1 RMSE / MAE
Based on the predicted registration matrices, the labeled points in all test OCTA images were reprojected to the corresponding test EMA images. Then, the root-mean-square-error (RMSE) [42] between the reprojected points and the labeled points was computed. In addition, the average maximum-absolute-error (MAE) [27], defined as the maximum reprojection error for each image pair, was also reported.
@@ 5.4.2 Success rate
The success rate was defined as the number of image pairs with successful registration over the total number of test image pairs. Since the definition of success varies, we defined a registration as successful when its RMSE $<$ 10 pixels or its MAE $<$ 10 pixels, based on the clinical tolerance.
5.4.3 Soft Dice / Masked Soft Dice
Dice [49] is widely used to evaluate the registration quality by calculating the pixel alignment between the warped source vessel masks and the target vessel masks. Soft Dice [27] has been proposed as an extension of Dice for assessing registration quality. For Soft Dice, CLAHE [69] was first applied to enhance the contrast of two input images and the Frangi vesselness filter [70] was then used to generate the vesselness probability masks of the two input images. We calculated Soft Dice by
$Soft\ Dice=\frac{2\sum_{i=1}^{N}min(F_{i}^{s^{\prime}},\ F_{i}^{t})}{\sum_{i=1}^{N}F_{i}^{s^{\prime}}+\sum_{i=1}^{N}F_{i}^{t}},$ (8)
@@ 5.4.3 Soft Dice / Masked Soft Dice
Figure 8: The average (a) Soft Dice and (b) Masked Soft Dice values over image pairs in MEMO by adding different x and y shifts to the ground truth registration. The top-left value in each figure represents the average Soft Dice or Masked Soft Dice value obtained by ground truth registration. All values are color-coded.
Additionally, we found that Soft Dice could not accurately represent performance when a relatively large VD difference was present between the two modalities, such as in our MEMO dataset. Moreover, it was particularly unreliable when the results of each competing method had relatively small differences. As most vessels in OCTA images do not exist in the corresponding EMA images, calculating Soft Dice based on every pixel is not ideal. Hence, when evaluating on the MEMO dataset, we extended Soft Dice to Masked Soft Dice which considered only the pixels within the ground truth segmentation masks of EMA images. Masked Soft Dice is defined as
$Masked\ Soft\ Dice=\frac{2\sum_{i=1}^{N}min(M_{i}^{e^{\prime}}F_{i}^{e^{\prime}},\ M_{i}^{e^{\prime}}F_{i}^{o})}{\sum_{i=1}^{N}M_{i}^{e^{\prime}}F_{i}^{e^{\prime}}+\sum_{i=1}^{N}M_{i}^{e^{\prime}}F_{i}^{o}},$ (9)
where $F^{e^{\prime}}$ and $F^{o}$ are the vesselness probability masks of the warped EMA images and the OCTA images, and $M^{e^{\prime}}$ represents the warped ground truth segmentation masks of EMA images. In Fig. 8, we demonstrated that Masked Soft Dice has better ability to assess the registration performance on our MEMO dataset, as it is less sensitive to the VD difference and noise.
@@ 6.1 The CF-FA Dataset
Methods
Success rate
(RMSE < 10)
Success rate
(MAE < 10)
RMSE MAE Soft Dice
Before Registration 0.00% 0.00% 68.35 77.03 -
SURF-PIIFD-RPM [43] 82.76% 75.86% 15.21 20.43 0.51
LoFTR [63] 41.28% 27.59% 45.13 52.29 0.51
SG [64] 82.76% 79.31% 23.68 31.40 0.52
@@ Methods
Content-Adaptive [28] 89.67% 75.86% 14.03 19.27 0.51
VDD-Reg (Ours) 100.00% 96.55% 3.19 5.54 0.51
The authors of Content-Adaptive [28] have demonstrated better performance on the CF-FA dataset by fine-tuning SuperPoint and adding an outlier rejection network.
@@ 6.2 The MEMO Dataset
Figure 9: Registration results of our method and the existing methods on a selected image pair from our MEMO dataset. The top row shows the grid images where the EMA and OCTA images are interlaced as small grids. The bottom row shows the overlay images of the EMA (green) and OCTA (orange) extracted vessels registered by each method. The vessel segmentation masks were generated by our method. The RMSE, MAE and Masked Soft Dice of each method are listed below each overlay image.
@@
Methods
Success rate
(RMSE < 10)
Success rate
(MAE < 10)
RMSE MAE Soft Dice Masked Soft Dice
Before Registration 6.67% 6.67% 88.48 98.17 - -
SURF-PIIFD-RPM [43] 6.67% 6.67% 58.48 100.13 0.42 0.41
LoFTR [63] 13.33% 13.33% 53.55 76.37 0.43 0.49
SG [64] 0.00% 0.00% 257.01 360.02 0.28 0.44
@@ Methods
Content-Adaptive [28] 33.33% 20.00% 78.69 105.03 0.41 0.48
VDD-Reg (Ours) 86.67% 66.67% 21.99 28.76 0.50 0.61
Table 4 shows the quantitative registration results of our method and the existing methods on our MEMO dataset. We also demonstrate the qualitative registration results of our method and the existing methods on a selected image pair from our MEMO dataset in Fig. 9. Due to the relatively large VD difference, all the methods performed worse compared to the results on the CF-FA dataset. Still, our VDD-Reg outperformed these five existing methods by a large margin (86.67%). From Table 4, we observed that SURF-PIIFD-RPM performed poorly (6.67%) on our MEMO dataset, which suggested that the hand-crafted features might be insufficiently powerful for this scenario. The two deep learning-based direct methods, LoFTR and SG, also produced unsatisfactory results (13.33% and 0%) due to the large distribution gap between their training dataset [72] and our MEMO dataset. CycleGAN-based with pretrained SuperPoint (SP) achieved relatively better performance compared to other competing methods, demonstrating its potential to solving difficult multimodal registration problems. Content-Adaptive, which is also a segmentation-based method, performed much worse (33.33%) than our method on the MEMO dataset. We attributed this to our use of annotated vessel segmentation masks from a single modality (EMA in our case). Different from our two-stage semi-supervised learning framework (LVD-Seg), Content-Adaptive trained the segmentation networks naively with style loss, the self-comparison loss and the image registration loss. To improve the segmentation quality, Content-Adaptive additionally guided the segmentation networks with mean phase images of input images using pixel-adaptive convolution (PAC) [51]. However, due to the high complexity of OCTA images, the OCTA mean phase images were usually too noisy to correctly guide the segmentation networks. On the other hand, our LVD-Seg framework used very few (e.g., three) annotated vessel segmentation masks from one modality to guide the segmentation networks to segment similar vessels in EMA and OCTA image pairs for both stages. In addition, the two-stage design also enhanced training stability when style loss was involved. These are particularly important for multimodal retinal image registration when a large VD difference exists between the two modalities.
@@ 7.1 Ablation Study on the Two-stage Learning Framework
Method Success rate RMSE Soft Dice Masked Soft Dice
@@
In this section, we investigated the benefits of each stage in the proposed two-stage semi-supervised learning framework (LVD-Seg) by removing one of the stages from the framework. The results are shown in Table 5. For Stage 1 only, we trained the EMA vessel segmentation network following the procedure described in Section 4.1.2 and used the same EMA vessel segmentation network for segmenting OCTA images. For Stage 2 only, we trained the segmentation networks with style loss only. In general, the performance of both variants decreased significantly. Stage 1 only performed the worst due to the limited number (three in our case) of annotated vessel segmentation masks used for supervised training, making the segmentation network generalize poorly on the test images and resulting in poor registration performance. Furthermore, Stage 1 only directly applied the trained EMA segmentation network on OCTA images, which did not work due to the relatively large VD difference between the two modalities. Although Stage 2 only worked better than Stage 1 only, it still significantly lagged behind our LVD-Seg. This implied that relying purely on style loss could let the training become unstable and resulted in unreliable registration. All these results emphasized the effectiveness of the proposed two-stage semi-supervised learning framework (LVD-Seg).
@@ 7.2 Ablation Study on Number of Required Vessel Masks
# of annotated masks Success rate RMSE Soft Dice Masked Soft Dice
@@ 7.3 Ablation Study on Data Used for Supervised Training
Datasets Success rate RMSE Soft Dice Masked Soft Dice
@@
7.4 Generalization Ability of the Proposed Method across Different NHPs
Table 8: Results of splitting the dataset by NHPs (Best results are marked in bold)
Methods Success rate RMSE Soft Dice Masked Soft Dice
[…]
VDD-Reg (Ours) 90.90% 9.38 0.52 0.62
So far, all experiments conducted on the MEMO dataset have utilized a train-test split that included both NHPs in each set. In this section, we further explored whether the proposed method trained on images of one NHP could be generalized to another. To achieve this, we divided the MEMO dataset by NHPs. Specifically, eight image pairs captured from one NHP were used as the training set, while the remaining 22 image pairs captured from the other NHP were used as the test set. Other than that, the default settings were adopted for training. The results are shown in Table 8. Given that CycelGAN-based achieved the second-best performance in Table 4, we compared our method with it in this study. Interestingly, we found that whether the training and test set included the same NHP did not affect the performance of our method, achieving a success rate of 90.9%. Additionally, our method outperformed CycelGAN-based by 9% in success rate, indicating its potential to register image pairs from different subjects.
@@ 8 Conclusion
In this paper, we present MEMO, the first public multimodal EMA and OCTA retinal image dataset. MEMO provides registration ground truth, EMA image sequences and OCTA projection images, desirable for various research fields. With MEMO, we first uncover a unique challenge of multimodal retinal image registration between modalities with large VD differences. After that, we propose a segmentation-based deep-learning registration framework, VDD-Reg, to deal with the large vessel density difference between EMA and OCTA in multimodal retinal image registration. Moreover, to train the segmentation module in our VDD-Reg, we design a novel two-stage semi-supervised learning framework, LVD-Seg, which combines supervised and unsupervised losses. Both quantitative and qualitative results demonstrate that VDD-Reg outperforms the existing methods in both small VD differences (i.e., CF-FA) and large VD differences (i.e., MEMO). Additionally, VDD-Reg requires as few as three annotated vessel segmentation masks to maintain its performance, which demonstrates its promising potential for registering other modalities.
Funding
The authors would like to acknowledge the funding support from the University of Maryland Strategic Partnership: MPowering the State, a formal collaboration between the University of Maryland College Park and the University of Maryland Baltimore Medical School. This work is also supported by NIH R01EY031731.
Disclosures
The authors declare no conflicts of interest.
Data availability
The MEMO dataset and CF-FA dataset underlying the results presented in this paper are available at https://chiaoyiwang0424.github.io/MEMO/ and [30].